OpenStudy (anonymous):

A pharmacist must mix 12 fl oz of cough syrup that contains 25% active ingredient with flavored syrup that contains no active ingredient. How many fluid oz of flavored syrup must the pharmacist add to create a mixture containing 10% active ingredient?

OpenStudy (anonymous):

Find the amount of active ingredient: 12 * 25% = 3 oz. active ingredient. Now you want to make 10% active ingredient with those 3 oz. 3/x = 10/100 x = 30 oz. //total volume So you need to add 18 oz to get 30 oz total.

OpenStudy (anonymous):

Thank you.